# Community-Cross-Instruct: Unsupervised Instruction Generation for Aligning Large Language Models to Online Communities

### v1

@@ Abstract
Social scientists use surveys to probe the opinions and beliefs of populations, but these methods are slow, costly, and prone to biases. Recent advances in large language models (LLMs) enable creating computational representations or “digital twins” of populations that generate human-like responses mimicking the population’s language, styles, and attitudes. We introduce Community-Cross-Instruct, an unsupervised framework for aligning LLMs to online communities to elicit their beliefs. Given a corpus of a community’s online discussions, Community-Cross-Instruct automatically generates instruction-output pairs by an advanced LLM to (1) finetune an foundational LLM to faithfully represent that community, and (2) evaluate the alignment of the finetuned model to the community. We demonstrate the method’s utility in accurately representing political and fitness communities on Reddit. Unlike prior methods requiring human-authored instructions, Community-Cross-Instruct generates instructions in a fully unsupervised manner, enhancing scalability and generalization across domains. This work enables cost-effective and automated surveying of diverse online communities.
@@
Zihao He, Rebecca Dorn, Siyi Guo, Minh Duc Chu, Kristina Lerman USC Information Sciences Institute {zihaoh, rdorn, siyiguo, mhchu}@usc.edu, lerman@isi.edu
@@ 1 Introduction
Social scientists use surveys and focus groups to learn the opinions, needs and concerns of diverse populations. However, designing surveys and recruiting participants is a slow and costly process, limiting the utility of these instruments for probing public opinion. Surveys are prone to biases, such as the social desirability bias Gordon (1987), where respondents may alter their responses to sensitive questions to appear more socially acceptable Bergen and Labonté (2020), non-response bias Hill et al. (1997), where participants fail to answer questions, and self-selection bias due to the choices participants make to participate in the survey Heckman (1990). In addition, social stigmas may taint responses Goel and Salganik (2010), especially for hard-to-reach and marginalized groups.
@@
Figure 1: Illustration of Community-Cross-Instruct to align an LLM to a community. (1) Open-ended instructions and multi-choice survey questions are generated by an advanced LLM from the community data. (2) A foundational LLM is aligned to community through instruction-tuning on the open-ended instructions. (3) The alignment of the finetuned LLM to the community is measured on the survey questions.
@@
Recent breakthroughs in generative AI and especially large language models (LLMs) enable new capabilities for creating computational representations of human populations — their digital twins El Saddik (2018) — by ingesting vast textual data they create, for example, in online discussion forums. These LLM-based models generate human-like responses that mimic the language, communication styles, and attitudes of populations they are aligned to, allowing us to probe their worldviews, biases, and sentiments in a cost-effective and automated manner. Previous works have leveraged such LLM-based representations to mine opinions and learn political attitudes of online communities Jiang et al. (2020); He et al. (2024b). However, they finetune GPT-2 Radford et al. (2019) directly on textual data from communities, where the finetuned model is not able to answer questions that may be in different format, e.g., multiple choice, as it is not an instruction-tuned model.
To address these challenges, we introduce Community-Cross-Instruct: a fully unsupervised framework for aligning LLMs to online communities through instruction-tuning. The finetuned LLMs serve as digital twins of communities, which can be automatically surveyed to elicit their views (Figure 1).
@@
Specifically, given a corpus of comments and submissions in different forums on Reddit, Community-Cross-Instruct uses an advanced LLM (Claude-3) to automatically curate (1) CommInst: a set of open-ended instructions (Fig. 2(a)) for community-specific instruction tuning, and (2) CommSurvey, a set of multi-choice survey questions (Fig. 2(b)). Each instruction in CommInst and each survey question in CommSurvey are paired with answers from different communities. We finetune foundational LLMs (GPT-3.5 or Llama-3) on CommInst, in order to align them to different communities, and evaluate the finetuned LLMs on CommSurvey to measure alignment. See Figure 3 for a high-level overview of the framework. Existing methods for self-supervised instruction generation rely on a seed set of instructions curated by domain experts Wang et al. (2022); Chen et al. (2024b), which is not generalizable. We overcome this bottleneck by incorporating the readily available community data into the instruction generation pipeline. Our Community-Cross-Instruct framework distills knowledge from communities using an advanced LLM to generate instructions for aligning foundational LLMs.
@@
  • We introduce Community-Cross-Instruct, a novel unsupervised framework for aligning foundational LLMs to online communities, by finetuning them on the automatically curated set of open-ended community-specific instruction-answer pairs (CommInst). The models’ alignment to communities is measured using another set of automatically generated multiple choice questions and answers (CommSurvey).
  • Using data from Reddit forums, we show that our method improves the fidelity of community representation (alignment) in two domains—politics and fitness—yielding significant alignment improvement over standard persona adaptation methods.
@@
Clarification on contributions.
We aim to align LLMs to online communities through instruction-tuning, so that the aligned LLMs can be flexibly surveyed to probe the communities’ views on different topics. To prepare data for instruction tuning, we propose Community-Cross-Instruct, and the automatic creation of the open-ended instruction-answer pairs (CommInst) is the core contribution. However, in order to demonstrate that the finetuned LLMs are indeed aligned to communities and can be treated as their digital twins, we propose to automatically generate survey questions and answers (CommSurvey) to efficiently evaluate LLM alignment. This is a secondary contribution of this paper, and we do not claim that CommSurvey is the best way to measure alignment. Therefore, our main goal is to propose a method to prepare LLMs as a proxy for surveying different online communities, rather than automatically creating such surveys.
On the use of the term alignment.
@@ 2 Problem Definition
A topical domain (e.g., politics or fitness) includes $n$ communities {$C_{1}$, $C_{2}$, …., $C_{n}$} each with different views and beliefs. Members of each community $C_{i}$ collectively author text corpus $D_{i}$ (e.g., discussions on Reddit forums) expressing views and exhibiting behaviors. Our goal is to align an LLM $f$ to each community $C_{i}$ using its texts $D_{i}$, such that the aligned LLM $f^{\prime}_{i}$ learns the complex mindset of the community and responds to inputs in the community’s voice. By administering surveys to the aligned LLMs $\{f^{\prime}_{1},f^{\prime}_{2},...,f^{\prime}_{n}\}$, we obtain responses from different communities, thereby capturing their ideological differences.
@@
To align an LLM $f$ to a community $C$, we finetune it on a set of demonstrations (instruction-answer pairs) Wang et al. (2023); Ouyang et al. (2022); Chen et al. (2024b) $I=\{(X_{j},Y_{j})\}_{j=1}^{m}$, where the instructions are open-ended questions probing the community’s views on different topics, and the corresponding answers are aligned with each community’s ideology. Figure 2(a) shows an example demonstration in the politics domain. We propose Community-Cross-Instruct (Fig. 3), a framework to automatically generate community-specific demonstrations $I$ with an advanced LLM $\hat{f}$ (Claude-3) based on the community’s text corpus $D$ and use these demonstrations to instruction-tune a foundational LLM $f$ (GPT-3.5 or Llama-3) through a process we call “Cross-Instruct”.
@@
Figure 2: Example of (a) an instruction from CommInst-Pol and (b) a survey question from CommSurvey-Pol. The instruction/survey question is paired with answers from different communities.
@@
Figure 3: Overview of Community-Cross-Instruct, with an illustrative example of the politics domain. (1) Data collection for each community within the desired domain. (2) BERTopic clusters the data and identifies prominent topics. (3) For each topic, the advanced LLM is prompted with (i) on-topic examples from each community and (ii) instruction-answer task definition and explanation. (4) Creation of (a) open-ended instruction-answer pairs and (b) multi-choice survey questions and answers, for fine-tuning and evaluating a sub-population representation foundational LLM respectively.
3 CommInst: A Collection of Community-Specific Instructions
To align LLMs to different communities within each domain, we curate a dataset CommInst for community-specific instruction tuning. CommInst-Pol and CommInst-Fit are the two subsets for the politics and fitness domains with 6,687 and 4,395 unique instructions respectively. Each instruction is an opinion-eliciting question about a topic relevant to the domain (Figure 2(a)), which appears in multiple demonstrations. Each demonstration’s instruction is paired with an open-ended answer from a different community.
3.1 Online Community Corpora Collection
@@
Reddit is a vibrant social media platform hosting discussion forums (subreddits) on a wide range of topics Hofmann et al. (2022); Chen et al. (2024a). In this paper we focus on two domains: politics and fitness, which contain distinctive subreddits, or communities, with complex social dynamics. Political subreddits are valuable expressions of diverse public opinions and viewpoints. In the age of rising polarization, LLMs can help researchers track the complex evolving ideological landscape and effectively elicit public opinions. Meanwhile, the fitness subreddits feature a wealth of sensitive, health-related conversations, and the latest diet and fitness fads, often discussed using inscrutable insider jargon. We can track the direction of these discussions, identifying emerging trends, risks, and potential health misinformation. This insight allows public health officials to promptly address harmful health advice and intervene where necessary. Additionally, it provides an early warning system to detect the spread of dangerous health practices, ensuring that corrective measures can be implemented swiftly to protect community health.
@@
We identify a set of representative subreddits for each domain based on personal knowledge and by querying ChatGPT. We manually aggregate and filter results, obtaining the following political online communities: r/Anarchism, r/Anarcho_Capitalism, r/Conservative, r/Libertarian, r/Republican, and r/democrats; For fitness, we investigate: r/Fitness, r/loseit, r/keto, r/xxfitness, and r/GettingShredded. More details of these communities are presented in Appendix A.1.
We collect comments and submissions from January 2019 to December 2022 and remove those that were deleted by the author or the moderator. The statistics of the collected comments and submissions are shown in Appendix A.1. To balance the number of comments and submissions across different communities, for each community, we keep a maximum of 140K comments and 30K submissions. Therefore, there are at most 170K texts, which makes up the text corpus $D$, including comments and submissions, for each community.
@@ 3.2 Topic Modeling
For a domain, denote the combined corpus from all $n$ communities by $\mathbb{D}=D_{1}\cup D_{2}\cup....\cup D_{n}$. We use BERTopic Grootendorst (2022) on $\mathbb{D}$ to identify topics $g$. More details about BERTopic can be found in Appendix B.1. After topic modeling, each text $d$ is assigned a topic $g(d)$. The number of topics detected in politics and fitness are 425 and 628 respectively. The top 10 most frequent topics in politics and fitness are shown in Appendix B.1 respectively. Politics topics include abortion, climate, guns, and racism. Fitness topics include body measures, keto diet, and weight lifting.
For each community $C_{i}$, we split the text corpus $D_{i}$ into smaller chunks $D_{i}=\{\hat{D}^{i}_{1},\hat{D}^{i}_{2},....\}$, where each chunk $\hat{D}$ consists of texts belonging to a single topic $g(\hat{D})$.
3.3 Open-ended Instruction-Answer Pair Generation
@@
For a specific domain, starting with an empty demonstration pool, we randomly sample without replacement a text chunk $D_{i}$ for each topic $t_{k}\in T$ from each community. We include the sampled text chunks on topic $t_{k}$ from the $n$ communities into a prompt and then use the prompt to query the advanced LLM to generate open-ended instructions and answers on the topic. The prompting template is shown in Appendix B.2. We query claude-3-opus-20240229 using the Anthropic API. In the prompt, we first specify the topic by the topic keywords, and then the comments in the sampled chunks from different communities. Next we ask the model to generate instructions that can be answered based on the texts, and that the instructions should elicit different responses from the $n$ communities. Each generated instruction is paired with different answers from different communities (Figure 2(a)). It is then added to the pool as $n$ demonstrations. Note that some topics may be discussed more in certain communities than others, so some instructions may be paired with answers from less than $n$ communities. We iteratively repeat this process until there are no text chunks left for topic $t_{k}$ from any community, until we shift to the next topic $t_{k+1}$. After the iterative generation process, the instruction-answer pool will contain demonstrations covering various topics, which will use to finetune foundational LLMs. Please refer to Appendix A.2 for more details about CommInst.
For fair evaluating the finetuned LLMs, we split CommInst-Pol in two ways:
  • Random split. CommInst-Pol is randomly split into training and eval sets: CommInst-Pol-train and CommInst-Pol-eval, in a ratio of 85:15.
  • Topical split. We randomly select 10 topics in politics and collect demonstrations related to the topics into CommInst-Pol-eval-topic. The remaining demonstrations are collected as CommInst-Pol-train-topic. In this way, CommInst-Pol-train-topic and CommInst-Pol-eval-topic has no topic overlap.
We repeat the same process for CommInst-Fit.
4 CommSurvey: A Collection of Community-Specific Survey Questions

### v2

@@ Abstract
Social scientists use surveys to probe the opinions and beliefs of populations, but these methods are slow, costly, and prone to biases. Recent advances in large language models (LLMs) enable the creation of computational representations or “digital twins” of populations that generate human-like responses mimicking the population’s language, styles, and attitudes. We introduce Community-Cross-Instruct, an unsupervised framework for aligning LLMs to online communities to elicit their beliefs. Given a corpus of a community’s online discussions, Community-Cross-Instruct automatically generates instruction-output pairs by an advanced LLM to (1) finetune a foundational LLM to faithfully represent that community, and (2) evaluate the alignment of the finetuned model to the community. We demonstrate the method’s utility in accurately representing political and diet communities on Reddit. Unlike prior methods requiring human-authored instructions, Community-Cross-Instruct generates instructions in a fully unsupervised manner, enhancing scalability and generalization across domains. This work enables cost-effective and automated surveying of diverse online communities [^1] Code and data are available at https://github.com/zihaohe123/community-cross-instruct.
@@
Zihao He, Minh Duc Chu, Rebecca Dorn, Siyi Guo, Kristina Lerman USC Information Sciences Institute {zihaoh, mhchu, rdorn, siyiguo}@usc.edu, lerman@isi.edu
**footnotetext: Equal Contribution.
@@ 1 Introduction
Social scientists use surveys and focus groups to learn the opinions, needs, and concerns of diverse populations. However, designing surveys and recruiting participants is a slow and costly process, limiting the utility of these instruments for probing public opinion. Surveys are prone to biases, such as the social desirability bias Gordon (1987), where respondents may alter their responses to sensitive questions to appear more socially acceptable Bergen and Labonté (2020), non-response bias Hill et al. (1997), where participants fail to answer questions, and self-selection bias due to the choices participants make to participate in the survey Heckman (1990). In addition, social stigmas may taint responses Goel and Salganik (2010), especially for hard-to-reach and marginalized groups.
@@
Figure 1: Illustration of Community-Cross-Instruct to align an LLM to a community. (1) Open-ended instructions and multi-choice survey questions are generated by an advanced LLM from the community data. (2) A foundational LLM is aligned to the community through instruction-tuning on the open-ended instructions. (3) The alignment of the finetuned LLM to the community is measured using the generated survey questions.
@@
Recent breakthroughs in generative AI and especially large language models (LLMs) enable new capabilities for creating computational representations of human populations — their digital twins El Saddik (2018) — by ingesting vast textual data they create, for example, in online discussion forums. These LLM-based models generate human-like responses that mimic the language, communication styles, and attitudes of populations they are aligned to, allowing us to probe their worldviews, biases, and sentiments in a cost-effective and automated manner. Previous works have leveraged such LLM-based representations to mine opinions and learn political attitudes of online communities Jiang et al. (2020); He et al. (2024b). However, these studies typically finetune models like GPT-2 Radford et al. (2019) directly on raw textual data from the communities, resulting in models that can perform text continuation but are limited in their ability to answer questions in structured formats, such as multiple choice. Moreover, raw community data often contains noise, irrelevant information, or off-topic discussions that can degrade model performance if not properly handled.
To address these challenges, we aim to align LLMs to communities by finetuning the models with high-quality instructions where the community perspectives are embedded. However, curating high-quality instructional data is a non-trivial process. Existing methods for self-supervised instruction generation rely on a seed set of instructions curated by domain experts Wang et al. (2022); Chen et al. (2024b), which is limited by the generalizability to a new domain. In this paper, we introduce Community-Cross-Instruct: a fully unsupervised framework for aligning LLMs to online communities through instruction-tuning. It incorporates the readily available community data into the instruction generation pipeline, which no human supervision is needed. Community-Cross-Instruct uses advanced LLMs to generate instruction-response pairs that better capture community perspectives in more structured and useful formats. These instructions-responses pairs are used to finetune foundational LLMs. The finetuned LLMs serve as digital twins of communities, which can be automatically surveyed to elicit their views. We show a high-level overview of our framework in Figure 1.
@@
Specifically, given a corpus of comments and submissions in different forums on Reddit, Community-Cross-Instruct uses an advanced LLM (GPT-4o) to automatically curate two instructional datasets: (1) CommInst: a set of open-ended instructions for community-specific instruction tuning, and (2) CommSurvey, a set of multi-choice survey questions for community-specific survey completion. Each instruction in CommInst and each survey question in CommSurvey are paired with responses from different communities (Figure 2). We finetune foundational LLMs (GPT-3.5 or Llama-3.1) on CommInst, in order to align them to different communities and evaluate the finetuned LLMs on CommSurvey to measure alignment.
@@
  • We introduce Community-Cross-Instruct, a novel unsupervised framework for aligning foundational LLMs to online communities, by finetuning them on the automatically curated set of open-ended community-specific instruction-response pairs (CommInst). The models’ alignment to communities is measured using another set of automatically generated multiple choice questions and answers (CommSurvey).
  • Using data from Reddit forums, we show that our method improves the fidelity of community representation (alignment) in two domains—politics and diet—yielding significant alignment improvement over standard persona adaptation methods.
@@
Using the term alignment.
@@ 2 Problem Definition
A topical domain (e.g., politics or diet) includes $n$ communities {$C_{1}$, $C_{2}$, …., $C_{n}$} each with different views and beliefs. Members of each community $C_{i}$ collectively author text corpus $D_{i}$ (e.g., discussions on Reddit forums) expressing views and exhibiting behaviors. Our goal is to align an LLM $f$ to each community $C_{i}$ using its texts $D_{i}$, such that the aligned LLM $f^{\prime}_{i}$ learns the complex mindset of the community and responds to inputs in the community’s voice. By administering surveys to the aligned LLMs $\{f^{\prime}_{1},f^{\prime}_{2},...,f^{\prime}_{n}\}$, we obtain responses from different communities, thereby capturing their ideological differences.
@@
To align an LLM $f$ to a community $C$, we finetune it on a set of demonstrations (instruction-response pairs) Wang et al. (2023); Ouyang et al. (2022); Chen et al. (2024b) $I=\{(X_{j},Y_{j})\}$, where the instructions are open-ended questions probing the community’s views on different topics, and the corresponding responses are aligned with each community’s ideology. Figure 2(a) shows an example demonstration in the politics domain. We propose Community-Cross-Instruct (Figure 3), a framework to automatically generate community-specific demonstrations $I$ with an advanced LLM $\hat{f}$ (GPT-4o) based on the community’s text corpus $D$ and use these demonstrations to instruction-tune a foundational LLM $f$ (GPT-3.5 or Llama-3.1) through a process we call “Cross-Instruct”.
@@
Figure 2: Example of (a) an instruction from CommInst and (b) a survey question from CommSurvey in the politics domain on the topic of marijuana. The open-ended instruction and survey question are paired with answers from different communities.
@@
Figure 3: Overview of Community-Cross-Instruct, with an illustrative example of the politics domain. (1) Data is collected for each community within the desired domain. (2) BERTopic clusters the data and identifies prominent topics. A chunk is a set of documents from a community on the same topic. Chunki,j represents the chunk from community $i$ on topic $j$. (3) For each topic, the advanced LLM is prompted with (i) on-topic chunks from each community and (ii) task definition of the instructional data generation (see Appendix B.2), which leads the LLM to generate (a) open-ended instruction-response pairs and (b) multi-choice question-answer pairs. Ri,k represents the response of community $i$ to instruction $k$; Ai,k represents the answer of community $i$ to question $k$. (4) The open-ended instructions across all topics, along with the corresponding responses of community $i$, are added to CommInsti, which is used to finetune a foundational LLM, to align the LLM to the community. (5) The multi-choice questions across all topics, along with the corresponding answers from community $i$, are added to CommSurveyi, which is used to evaluate the finetuned LLM.
3 Online Community Corpora Collection
@@
Reddit is a vibrant social media platform hosting discussion forums (subreddits) on a wide range of topics Hofmann et al. (2022); Chen et al. (2024a). In this paper, we focus on two domains: politics and diet, which contain distinctive subreddits, or communities, with complex social dynamics. Political subreddits are valuable expressions of diverse public opinions and viewpoints. In the age of rising polarization, LLMs can help researchers track the complex evolving ideological landscape and effectively elicit public opinions. Meanwhile, the diet subreddits feature a wealth of sensitive, health-related conversations, and the latest diet and diet fads, often discussed using inscrutable insider jargon. We can track the direction of these discussions, identifying emerging trends, risks, and potential health misinformation. This insight allows public health officials to promptly address harmful health advice and intervene where necessary. Additionally, it provides an early warning system to detect the spread of dangerous health practices, ensuring that corrective measures can be implemented swiftly to protect community health.
@@
We identify a set of representative subreddits for each domain based on personal knowledge and by querying ChatGPT. We manually aggregate and filter results, obtaining the following five political online communities: r/Liberal, r/NeutralPolitics, r/Anarcho_Capitalism, r/Conservative, and r/AskThe_Donald; For diet, we investigate three communities: r/keto, r/WeightLossAdvice, and r/EDAnonymous. More details of these communities are presented in Appendix A.1.
We collect comments and submissions from January 2019 to December 2022 and remove those that were deleted by the author or the moderator through PushShift [^2] https://pushshift.io. The statistics of the collected comments and submissions are shown in Appendix A.1. We do not differentiate between comments and submissions, and treat each comment and submission as a separate document. In subsequent sections of this paper, we use “comment” to refer to both “comment” and “submission”.
4 Instructional Data Generation
For each domain, we curate a collection of (1) instructional datasets CommInst for community-specific instruction tuning, which is used to finetune LLMs to views of different communities, and (2) survey datasets CommSurvey for evaluating the alignment of the finetuned LLMs’ views to the communities using survey questions. The pipeline is depicted in Figure 3.
@@ 4.1 Topic Modeling
For a domain, denote the combined corpus from all $n$ communities by $\mathbb{D}=D_{1}\cup D_{2}\cup....\cup D_{n}$. We use BERTopic Grootendorst (2022) on $\mathbb{D}$ to identify topics $T$. More details about BERTopic can be found in Appendix B.1. After topic modeling, each text $d$ is assigned a topic $t(d)$. For each community $C_{i}$, we split the text corpus $D_{i}$ into smaller chunks $D_{i}=\{\hat{D}^{i}_{1},\hat{D}^{i}_{2},....\}$, where each chunk $\hat{D}^{i}$ consists of 50 randomly sampled texts belonging to a single topic $t(\hat{D}^{i})$. A chunk $\hat{D}^{i}$ on topic $t(\hat{D}^{i})$ is considered an occurrence of the topic. To investigate different views across the communities, we only keep the topics that are discussed by at least $n-1$ communities. To balance the occurrences of different topics, for each topic in a community, we keep a maximum of 5 chunks. As a result, for politics and diet, there are 41 and 148 topics respectively. We present the keywords of 10 topics that appear 5 times in all communities in Appendix B.1. Politics topics include COVID vaccines, abortion, climate, guns. diet topics include body measures, keto diet, and eating disorder recovery.
4.2 Open-ended Instructions Generation
@@
For a specific domain with $n$ communities, we initialize $n$ empty demonstration pools $\{\textsc{CommInst}_{i}\}_{i=1}^{n}$. For a topic $t_{k}\in T$ that is discussed by $m$ communities [^3] As we require a topic to appear at least $n-1$ communities, $m=n$ or $m=n-1$., from each community $C_{i}$, we randomly sample without replacement a text chunk $\hat{D}^{i}$ on $t_{k}$. We include the $m$ sampled text chunks $\{\hat{D}^{1},\hat{D}^{2},...,\hat{D}^{m}\}$ on topic $t_{k}$ into a prompt and then use the prompt to query the advanced LLM (GPT-4o) to generate open-ended instructions and responses on the topic. The prompting template is shown in Appendix B.2. In the prompt, we first specify the topic by the topic keywords, and then the comments in the sampled chunks from different communities. Next, we instruct the model to generate an instruction that can be answered based on the texts, and that the instruction should elicit different responses from the communities. In addition, we instruct the model to not rely on its pre-knowledge about the community, but solely focus on the given texts. The generated instruction $X$ is paired with different responses $\{Y_{i}\}_{i=1}^{m}$ from different communities, as shown in Figure 2(a). Each demonstration $\{X,Y_{i}\}$ is then added to the corresponding pool for the $i$-th community. We iteratively repeat this process until there are fewer than $n-1$ communities that have chunks left on topic $t_{k}$, then we shift to the next topic $t_{k+1}$. After the iterative generation process, each demonstration pool $\textsc{CommInst}_{i}$ is used to finetune a foundational LLM for community $C_{i}$. Please refer to Appendix B.3 for more details about CommInst.
4.3 Multi-choice Surveys Generation